Over a time interval of 1.61 years, the velocity of a planet orbiting a distant star reverses direction, changing from +20.1 km/s to -20.7 km/s. Find (a) the total change in the planet's velocity (in m/s) and (b) its average acceleration (in m/s2) during this interval. Include the correct algebraic sign with your answers to convey the directions of the velocity and the acceleration.
Answer #1

vi=+20.1km/s, vf=-20.7km/s, Δt=1.61yers = (1.61*265*24*60*60)s = 3.686*10^7 s

a)Δv = (vf - vi) = (-20.7km/s -20.1km/s) = -40.8 km/s = -40.8*1000 m/s = -40800 m/s

b)vave = Δv/Δt = (-40800)/(3.686*10^7) = -1.11*10^-3 m/s2